You are using water to dilute small amounts of chemicals in the laboratory, drop by drop. How many drops of water are there in a 1.0L bottle ?
1
Expert's answer
2019-10-09T10:41:32-0400
"V_{drop} = 4/3\pi r^3"

"mg=2\pi r\gamma"

"m=V_{drop} \rho"

Where V drop - volume of one drop, m - mass one drop, r - radius of drop, rho - density of water, gamma - surface tension of water.

"4/3\pi r^3 \rho g= 2\pi r \gamma"

"r =( 3\gamma/2g\rho)^{0.5}"

"r=0.0033 m"

"V_{drop} = 0.15 ml"

"Quantity _{drops}= 1000\/0.15= 6542"